OpenStudy (anonymous):

Find the derivative in terms of x for 4sin(2y)cos(x)=2

OpenStudy (anonymous):

I'm really struggling on this if someone could please help

OpenStudy (kira_yamato):

\[4\sin(2y)\cos(x)=2 \implies \sin(2y) = \frac{1}{2\cos(x)} = \frac{1}{2}\sec(x)\] \[\implies 2y = \sin^{-1}\left( \frac{1}{2}\sec(x) \right)\] \[\implies y = \frac{1}{2}\sin^{-1}\left( \frac{1}{2}\sec(x) \right)\]

OpenStudy (kira_yamato):

You could use the chain rule from here.

OpenStudy (kira_yamato):

\[\frac{d}{dx}\left( \sin^{-1}x \right) = \frac{1}{\sqrt{1-x^2}}\] \[\frac{d}{dx}\left( \sec x \right) = \sec x \tan x\]